# Crosslingual Capabilities and Knowledge Barriers in Multilingual Large Language Models

### v1

@@ B.3 WikiText-103 subset: Harry Potter-related documents
Appendix C Additional experimental results
@@
Figure 7 in the main paper presents the Harry Potter Quiz evaluation results on Llama2-7B and Llama3-8B models fine-tuned on general knowledge corpora (i.e., WikiText-2). Figure 9 presents additional results for the Llama2-13B (left) and Mistral 7B (right) models. (1) The trends are consistent with those observed for Llama2-7B and Llama3-8B, where fine-tuning on a mixed-language general corpus, WikiText-2, enhances the models’ performance on the domain-specific HP Quiz task across multiple languages, including English. (2) Word-level language mixing is generally most effective for Llama2-13B, whereas sentence-level mixing is more effective for Mistral-7B.
@@
Figure 10: Performance of Llama2-7B models on MMLU variant benchmarks. Fine-tuning on mixed language WikiText-103 generally outperforms fine-tuning on English WikiText-103 or using the un-finetuned model.
@@
Figure 11: Performance of Llama3-8B models on MMLU variant benchmarks. Fine-tuning on mixed language WikiText-103 generally outperforms fine-tuning on English WikiText-103 or using the un-finetuned model.
@@
Table 4 in the main paper presents the English MMLU and mixup MMLU evaluation results on Llama2-7B and Llama3-8B models fine-tuned on general knowledge corpora (i.e., WikiText-103). Here we present additional results for Llama2-7B (Figure 10) and Llama3-8B (Figure 11) on more MMLU variant benchmarks, including full translation, question translation, options transition, and ground-truth option translation. We report the average accuracy (with *) across four non-English languages {fr, de, es, and it } for those settings.
(1) As shown in Figure 10 and Figure 11, models fine-tuned on mixed language WikiText-103 (whether at the word level or sentence level) generally achieve better performance than those fine-tuned on the original English WikiText-103 or the un-finetuned models, especially in the GT-option translated and mixup translated MMLU setups. These two evaluation setups originally had the lowest performance for the un-finetuned model, and thus the cross-lingual ability gains after fine-tuning are more apparent. These results suggest that multiple language switches during fine-tuning enable LLMs to better understand and process multilingual input and leverage cross-lingual knowledge for commonsense reasoning tasks. (2) An exception to this trend is observed with the GT-option translated MMLU under the 5-shot biased demonstrations setting for Llama3-8B, where performance drops. This drop is likely due to the un-finetuned Llama3-8B’s stronger tendency to follow biased demonstrations, using a shortcut to select the non-English option as the answer. (3) Fine-tuning models on a mixed-languages corpus performs better than other models across different 0-shot and few-shot scenarios, particularly in the 0-shot setting and 5-shot English demonstrations setting. While 5-shot biased demonstrations lead to the best performance, they are less applicable than English demonstrations in real-world scenarios, as we cannot know in advance the language mixing pattern of user queries.

### v2

@@ D.3 WikiText-103 subset: Harry Potter-related documents
Table 7: COMET scores for machine translation tasks evaluated on FloRes-101 benchmark using multilingual LLMs (Llama2-7B, Mistral-7B, Llama2-13B, Llama3-8B, GPT-3.5, and GPT-4), models trained on parallel corpora (NLLB-3.3B), and an industrial-grade translation API (Google Translate). Multilingual LLMs achieve competitive translation performance against dedicated translation models and the translation API.
Llama2-7B $81.67$ $84.54$ $84.76$ $85.17$ $84.04$ $87.61$ $87.96$ $85.60$ $86.47$ $86.91$
Llama3-8B $73.89$ $81.19$ $81.03$ $81.16$ $79.32$ $88.52$ $88.61$ $86.45$ $87.03$ $87.65$
Mistral-7B $76.18$ $78.46$ $80.04$ $76.64$ $77.83$ $87.73$ $88.05$ $85.72$ $86.11$ $86.90$
Qwen2.5-7B $85.40$ $87.35$ $86.15$ $86.56$ $86.37$ $88.89$ $89.24$ $86.78$ $88.17$ $88.27$
Llama-3.1-8B $80.19$ $85.73$ $84.94$ $84.90$ $83.94$ $88.89$ $88.32$ $85.64$ $86.72$ $87.39$
TowerBase-7B-v0.1 $85.43$ $88.01$ $86.10$ $88.44$ $86.99$ $88.13$ $88.41$ $85.16$ $87.74$ $87.36$
GPT-3.5 $87.53$ $86.97$ $86.40$ $86.46$ $86.84$ $89.14$ $89.42$ $87.47$ $88.41$ $88.61$
GPT4 $88.18$ $88.13$ $86.64$ $85.33$ $87.07$ $89.71$ $89.56$ $87.41$ $88.14$ $88.71$
NLLB-3.3B $87.33$ $87.44$ $86.88$ $88.26$ $87.48$ $79.65$ $87.44$ $85.64$ $84.13$ $84.72$
Google Translate $89.39$ $89.22$ $87.23$ $89.37$ $88.80$ $90.01$ $89.92$ $87.55$ $88.54$ $89.01$
Appendix E Additional experimental results
E.1 Crosslingual capabilities of LLMs
Machine translation
Tb. 7 report the COMET score on FLoRes-101 benchmark (Goyal et al., 2022) for two directions per language: en $\rightarrow$ X and X $\rightarrow$ en. It shows that multilingual LLMs;s translation ability is quite competitive when compared to translation models explicitly trained on parallel corpora or industrial-grade translation APIs.
Embedding analysis
As shown in Fig. 11, for the four LLMs multilingual, including Llama2-7B, Llama2-13B, Mistral-7B and Llama3-8B [^10] We focus primarily on open-source models due to the cost associated with querying embeddings from proprietary models., the embeddings of the original text and its mixed-translated counterpart exhibit a high degree of similarity, with their difference vectors clustering around the origin. This observation stands in stark contrast to the scenario where English words are replaced with random tokens. It implies the explicit crosslingual capabilities of the multilingual LLMs.
E.2 Crosslingual knowledge barriers in LLMs
Evaluation strategy
We follow prior work on LLM evaluation (Zheng et al., 2024) to use two evaluation strategies: (1) Open-source: access the output probabilities of option ID tokens A/B/C/D and predict argmax. (2) Closed-source: compare the golden answer with the 1st generated token (decoding temperature=0), as the logits are not available for most closed-source models.
[…]
Table 8: Comparing two evaluation strategies for open-source LLMs: option ID token with maximum probability and first new token.
Model Eval English MMLU Mixup MMLU
Max prob 1st token Max prob 1st token
Llama2-7B 0-shot $41.53$ $37.74$ $32.18$ $27.47$
5-shot $45.88$ $45.90$ $36.92$ $36.96$
Mistral-7B 0-shot $60.21$ $58.41$ $47.86$ $42.29$
5-shot $62.57$ $62.54$ $51.07$ $51.05$
Llama3-8B 0-shot $60.54$ $62.11$ $48.62$ $50.13$
5-shot $65.00$ $65.39$ $51.65$ $52.01$
Crosslingual evaluation of additional models on MMLU knowledge
We present additional results for Llama2-7B, Zamba-7B, and Mixtral-8x7B. Figure 12 shows the monolingual evaluation of LLMs on MMLU, fully translated for non-English languages. The models consistently achieve higher accuracy in English compared to other languages.
Figure 13 displays the accuracy of LLMs on MMLU variant benchmarks. We observe a significant drop in accuracy under crosslingual MCQ evaluation, especially for ground-truth translated MMLU variant, indicating cross-lingual knowledge barriers. The barrier is more pronounced in Llama2-7B and Zamba-7B than in Mixtral-8x7B, possibly due to the larger capacity and multilingual capabilities of Mixtral-8x7B.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: Monolingual evaluation of LLMs on MMLU (fully translated for non-English languages). LLMs consistently perform better at answering multi-choice questions in English than in other languages.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: Crosslingual evaluation of LLMs on MMLU variant benchmarks. The bars with * denotes the average accuracy across {fr, de, es, and it}. LLMs perform better at answering MCQs in English than in mixed-language settings, especially the ground truth option and mixup translation, indicating the existence of cross-lingual knowledge barriers. Due to budget constraints, GPT-4 is evaluated only in the most challenging settings.
E.3 Mixed-language fine-tuning
@@
Fig. 7 in the main paper presents the Harry Potter Quiz evaluation results on Llama2-7B and Llama3-8B models fine-tuned on general knowledge corpora (i.e., WikiText-2). Fig. 14 presents additional results for the Llama2-13B (left) and Mistral 7B (right) models. (1) The trends are consistent with those observed for Llama2-7B and Llama3-8B, where fine-tuning on a mixed-language general corpus, WikiText-2, enhances the models’ performance on the domain-specific HP Quiz task across multiple languages, including English. (2) Word-level language mixing is generally most effective for Llama2-13B, whereas sentence-level mixing is more effective for Mistral-7B.
@@
Figure 15: Performance of Llama2-7B models on MMLU variant benchmarks. Fine-tuning on mixed language WikiText-103 generally outperforms fine-tuning on English WikiText-103 or using the non-fine-tuned model.
@@
Figure 16: Performance of Llama3-8B models on MMLU variant benchmarks. Fine-tuning on mixed language WikiText-103 generally outperforms fine-tuning on English WikiText-103 or using the non-fine-tuned model.
Figure 17: After sentence-level mixed-lang FT, embeddings of original English text & mixed-language-translated text are more closely aligned, indicating a stronger knowledge correlation between En & other langs.
Figure 18: Fine-tuning on a mixed-language domain-specific corpus (i.e., Harry Potter related documents from WikiText-103) generally enhances the performance on the Harry Potter Quiz dataset across multiple languages, including English.
@@
Tb. 3 in the main paper presents the English MMLU and mixup MMLU evaluation results on Llama2-7B and Llama3-8B models fine-tuned on general knowledge corpora (i.e., WikiText-103). Here we present additional results for Llama2-7B (Fig. 15) and Llama3-8B (Fig. 16) on more MMLU variant benchmarks, including full translation, question translation, options transition, and ground-truth option translation. We report the average accuracy (with *) across four non-English languages {fr, de, es, and it } for those settings.
(1) As shown in Fig. 15 and Fig. 16, models fine-tuned on mixed language WikiText-103 (whether at the word level or sentence level) generally achieve better performance than those fine-tuned on the original English WikiText-103 or the non-fine-tuned models, especially in the GT-option translated and mixup translated MMLU setups. These two evaluation setups originally had the lowest performance for the non-fine-tuned model, and thus the cross-lingual ability gains after fine-tuning are more apparent. These results suggest that multiple language switches during fine-tuning enable LLMs to better understand and process multilingual input and leverage cross-lingual knowledge for commonsense reasoning tasks. (2) An exception to this trend is observed with the GT-option translated MMLU under the 5-shot biased demonstrations setting for Llama3-8B, where performance drops. This drop is likely due to the non-fine-tuned Llama3-8B’s stronger tendency to follow biased demonstrations, using a shortcut to select the non-English option as the answer. (3) Fine-tuning models on a mixed-languages corpus performs better than other models across different 0-shot and few-shot scenarios, particularly in the 0-shot setting and 5-shot English demonstrations setting. While 5-shot biased demonstrations lead to the best performance, they are less applicable than English demonstrations in real-world scenarios, as we cannot know in advance the language mixing pattern of user queries.